Find the wavelength (in nm) of the photon emitted when an electron in an H atom undergoes a transition from n=5 to n=2
1
Expert's answer
2016-05-10T11:04:03-0400
1/λ = R*(1/22 – 1/52), where R = 109737,3157 сm−1 (Rydberg constant)
1/λ = 109737.3157*0,21 =23044.8363 сm−1
λ = 1/23044.8363 сm−1 = 0.000043394 cm = 433.94 nm (434 nm)

Answer: The wavelength is 433.94 nm (434 nm)